# TOPA: Extend Large Language Models for Video Understanding via Text-Only Pre-Alignment

### v1

@@ 3.4 Implementation Details
We leverage Llama2-7B, Llama2-13B [51] and Llama3-8B as the LLM backbone. Additionally, we employ the Llama-adapter [74] with an adaptation embedding length of 50 for efficient finetuning. We utilize CLIP-ViT-L as the multimodal encoder. We employ a simple linear layer to project the CLIP feature into the LLM feature space. The CLIP model and LLM backbone are frozen. The projection layer and additional Llama-adapter are trainable. For text-only pre-alignment, we uniformly sample the Tideos into 10 frames. We train the model on a mixture of tasks comprising Tideo summarization, Tideo QA, multi-choice Tideo QA with the ratio of 1:1:2. TOPA-Llama2-7B and TOPA-Llama3-8B are trained on four 40G-A100 GPUs in one day. TOPA-Llama2-13B is trained in two days. For zero-shot inference, we construct a memory for cross-modal projection, consisting of 2M CLIP text features sampled from captions in the TextVid dataset. More training details of TOPA and baselines are included in Appendix C.2.
@@ 4 Experiments
TOPA enables the LLM to perform various video understanding tasks as shown in Figure 2. In this section, we evaluate TOPA on multi-choice video QA and video captioning tasks. Section 4.1 evaluates TOPA on NeXT-QA [62], STAR [61], TVQA [20], and recent challenging EgoSchema [34] benchmarks with the zero-shot setting. We further evaluate TOPA on multi-choice video QA with the finetuning setting (Section 4.2) and zero-shot video captioning task (Section 4.3). In Section 4.4, we conduct ablation study on the LLM prior and input video frames. We mainly compare TOPA against four categories of video understanding methods as follows:
@@
(2) Adapt image MLLMs for video understanding [18, 71]. These approaches aim to extend the image understanding capabilities of recent vision-language models (VLMs) to video understanding. Specifically, SeViLa [71] utilizes BLIP-2 for localizing and understanding key frames of a video. IG-VLM [18] converts video into a composite image by arranging the video frames into a grid layout.
(3) Instructed Video MLLMs [24, 58, 73, 78]. Recent research employ video instruction-tuning for video-LLM alignment to enhance video-language understanding.
(4) LLM-based video agents [37, 52, 53, 55, 72]. This line of work leverages LLMs like GPT-3.5 and GPT-4 as an agent to understand a video by designing and executing a series of actions. The language-only agents perceive visual information via recent foundation VLMs (e.g., CLIP [43], BLIP-2 [23], LaViLa [77] and PALI [8]).
(5) Our text-only pre-alignment. Different from the above works, TOPA leverages the proposed TextVid dataset for video-LLM pre-alignment, enabling the LLM to process continuous features. Thus, performing video understanding tasks.
@@ 4.1 Zero-Shot Evaluation on Multi-Choice Video QA
In Section 4.1.1, we compare TOPA with a wide range of approaches on a recent challenging long-form video understanding benchmark, i.e., EgoSchema [34]. In Section 4.1.2, we further evaluate TOPA on conventional multi-choice video QA benchmarks including NeXT-QA [62], STAR [61] and TVQA [20]. In Section 4.1.3, we evaluate TOPA on recent comprehensive video understanding benchmark, i.e., MVBench [25]. The details of these benchmarks are included in Appendix C.1.
@@ 4.1.1 Zero-shot Results on EgoSchema
denotes results on EgoSchema subset. Results of InterVideo and FrozenBiLM are sourced from  [34]. Results of SeViLA are sourced from [40].
@@
Human Eval [34][NeurIPS 2023] - - 75.0%
Gemini-1.5-Pro [44][arXiv 2024.2 ] - Gemini-1.5-Pro 63.2%
@@
FrozenBiLM [67][NeurIPS 2022] - - 26.9%
InterVideo [56][arXiv 2022.12] - - 32.1%
LongViViT [40][CVPR  2024 ] - - 33.3%
MC-ViT-L [4][ICML  2024 ] - - 44.4%
InterVideo2s3-6B [58][arXiv 2024.3 ] - - 41.1%
@@
SeViLA [71][NeurIPS 2023] BLIP-2 FLAN-T5-XL [10] 22.7%
IG-VLM [18][arXiv 2024.3 ] LLaVA v1.6-7B Vicuna-7B 35.8%
IG-VLM [18][arXiv 2024.3 ] LLaVA v1.6-13B Vicuna-13B 47.0%
@@
Vamos [53][arXiv 2023.11] BLIP-2 Llama2-13B 36.7%
Vamos [53][arXiv 2023.11] BLIP-2 GPT-3.5 41.2%
Vamos [53][arXiv 2023.11] BLIP-2 GPT-4 48.3%
LLoVi [72][arXiv 2024.3 ] LaViLa GPT-3.5 50.3%
VideoAgent [55][arXiv 2024.3 ] LaViLa GPT-4 54.1%
MoReVQA [37][arXiv 2024.4 ] PALI-3-5B [8] PaLM-2 [2] 51.7%
(Our Text-Only Pre-Alignment) TOPA CLIP-ViT-L Llama2-7B 41.2%
TOPA CLIP-ViT-L Llama3-8B 44.2%
TOPA CLIP-ViT-L Llama2-13B 51.0%
Results on Egoschema. Table 1 shows the results. We compare our method against a range of recent approaches in video understanding. Our proposed text-only pre-alignment framework, despite training without real videos, shows impressive results on the Egoschema benchmark. TOPA outperforms previous image-based adaptation approach IG-VLM and video agents LLoVi and Vamos with the same scale LLM (Llama2-7B and Llama2-13B). Moreover, TOPA-Llama2 demonstrates consistent improvements when scaled up with a larger LLM backbone, indicating the effectiveness of LLMs in complex video-language understanding scenarios.
Discussion 1: The necessity of high-quality language supervision for video understanding. Recent video pre-training approaches like LongViVit [40] and InternVideo [56], despite training on million-level web video-text data, show inferior performance on Egoschema evaluation. These results highlight the inefficacy and inefficiency of conventional contrastive pre-training in understanding long-form videos, primarily due to noisy and simplistic language supervision. In contrast, our TOPA, trained on 721K virtual video (i.e., Tideo) with high-quality language supervision, shows impressive results on Egoschema. It indicates that, unlike image understanding which significantly benefits from leveraging web language as supervision, video understanding may require more precise and accurate language supervision to better capture the complex visual dynamics.
Discussion 2: Video agents versus end-to-end video-LLM modeling. Video agents have shown impressive results on the EgoSchema benchmark, aided by advanced LLMs and VLMs. However, a significant limitation of these approaches is their heavy reliance on the powerful LLMs. For example, the accuracy of Vamos drops by -11.6% when the GPT-4 is replaced with Llama2-13B. The reliance on powerful closed-source LLMs restricts its application fields and introduces external overheads. Additionally, the inference speed of these approaches is another concern, since it involves multiple interactions with both VLMs and LLMs. In contrast, end-to-end video-LLM models, which condense the video into a sequence of embeddings as the input of LLM, are more efficient. Moreover, video agents make decisions based on the language format clues collected by VLMs. Converting the video content into language clues may lead to a limited upper bound compared to end-to-end modeling.
@@ 4.1.2 Zero-shot Results on NExT-QA, STAR and TVQA
4.1.3 Zero-shot Results on MVBench
MVBench [25] is a recent video-language understanding benchmark that covers 20 challenging video tasks, regrouped from existing video-language benchmarks. For details of these 20 video understanding tasks, please refer to  [25]. Table 3 shows the results. TOPA demonstrates impressive results compared to previous image MLLMs and video MLLMs. It excels particularly in tasks requiring high-level video-language understanding, such as Scene Transition (ST), Episodic Reasoning (ER), and Unexpected Action (UA). Surprisingly, TOPA also excels in the Action Localization (AL) task, which requires identifying the moment an action occurs. This indicates that the text-only pre-alignment enables the LLM to understand visual temporal sequences. However, TOPA struggles with tasks that demand fine-grained visual understanding, such as Moving Direction (MR), Action Antonym (AA), and Object Shuffle (OS). A common challenge in these tasks is the requirement for detailed visual understanding. For example, Action Antonym involves identifying the direction of an action, while Object Shuffle involves locating objects. TOPA struggles in these fine-grained visual tasks since it is trained with CLIP text features. The modality gap between CLIP text features and image features hinders TOPA from capturing visual details. Further video instruction tuning might address this limitation, which we leave for future work. We provide qualitative results in Figure 4 - 7 to illustrate TOPA’s advantages and limitations on various video understanding tasks.
@@ 4.1.3 Zero-shot Results on MVBench
Video MLLMs: All models take 16 frames as input.
@@
Table 3: Evaluation results on MVBench. The results of other approaches are sourced from  [25]. We gray out the results of VideoChat2 because it utilizes extensive manually annotated video data from downstream video benchmarks for video instruction tuning.
@@
4.2 Finetuning Results
In this section, we further finetune the pre-aligned TOPA models to study the benefits of TOPA for downstream supervised learning. During finetuning, TOPA directly takes the video feature as input without the cross-modal projection. More finetuning details for each dataset are provided in Appendix C.2. Table 4 shows the finetuning results on multi-choice video QA dataset including NExT-QA, STAR and TVQA. For comparison, we include baseline models without text-only pretraining. Our text-only pre-alignment consistently improves the performance across three benchmarks. Notably, TOPA-Llama2-7B achieves 67.1% accuracy on TVQA, outperforming other approaches by a large margin. Additionally, in the zero-shot evaluation (Table 2), TOPA-Llama2-7B shows inferior results on STAR compared to SeViLA, due to its limitation in handling fine-grained visual details. However, in this finetuning setting, TOPA-Llama2-7B outperforms SeViLA, indicating that finetuning on downstream video data effectively mitigates TOPA’s limitations in processing fine-grained visual details. These results suggest that our text-only pre-alignment, despite not training with real video data, has a similar effect to conventional video-text pre-training.
@@ 4.2 Finetuning Results
Data Efficient Finetuning Figure 3 shows the results of finetuning LLMs with various ratios of training data. TOPA trained with 10% data achieves 64.7% Top 1 accuracy on NeXT-QA benchmark, significantly outperforming the baseline that without text-only pre-alignment. Besides, when trained with less than 20% data, the baseline model even performs worse than TOPA-ZS on NeXT-QA and TVQA, demonstrating the effectiveness of TOPA in limited annotated data scenarios.
@@ 4.4 Ablations
5 Limitations.
Modality gap in CLIP. Despite the fact that TOPA achieves impressive results. A significant limitation in TOPA is the gap between the CLIP text feature and CLIP image feature. On the one hand, we use the CLIP text feature for pre-alignment, while inference is with the CLIP visual feature. The modality gap makes the performance degrades, despite we employ a modality projection mechanism to mitigate it. On the other hand, the CLIP text features cannot fully capture the fine-grained visual details present in actual images, such as object locations and relationships. This limitation causes TOPA to struggle in scenarios where questions involve detailed visual information, as shown in Appendix A.3.
6 Broader Impact
@@ 7 Conclusions
Appendix B: The details of proposed TextVid dataset.
Appendix C.1, The details of benchmarks.
Appendix C.2, The training details of TOPA.
Appendix D: The prompts used in this paper.
Appendix E: Examples from TextVid.
@@ A.1 Further Discussion on Multi-Choice Video QA Task
TOPA-LLama2-13B LLM Logits 67.5 41.6 -25.9
@@
The performance gap between the EgoSchema subset and full set. Previous work [40, 4] highlights a huge performance gap between the subset and the full set of EgoSchema as shown in Table 8. We observe that such a performance gap phenomenon also occurs in approaches based on LLM logits. However, but it diminishes or even disappears in methods employing LLM selection. We find that this phenomenon may be attributed to differences in the linguistic structures of the choices, as shown below. The choices in the subset often differ in several key works like “create”, “repair” and “clean”. The similarity or logit can effectively identify this keyword-level difference to select a more appropriate choice. Conversely, the choices in the full set display more substantial linguistic differences. These variations introduce significant language biases, i.e., some sentences naturally receive higher logits in LLM, complicating the reliance on similarity or logit for choice selection. In contrast, LLM selection methods take all the choices within the context, allowing the LLM to leverage its robust contextual understanding to select the correct choice.
@@ A.3 Qualitative Results and Analysis
We present qualitative results to illustrate the capabilities and limitations of TOPA across various video understanding tasks. Figure 8 shows qualitative results on the NExT-QA validation set. Figure 9 shows qualitative results on the EgoSchema subset. Figure 4 - 7 shows qualitative results on 20 video understanding tasks from MVBench.
@@
Appendix B The details of proposed TextVid dataset
We utilize Gemini Pro 1.0 API for our data generation process. We prompt the LLM to create textual videos along with associated annotations. To ensure a diverse dataset that covers a wide of domains, we add condition prompts including different themes, video captions, video events, and the names of main objects. Specifically, we leverage video titles from Howto100M [35], video captions from WebVid2M [3], video events from Ego4D [14], and object from Wordnet [36] as conditions to generate diverse textual videos. For Ego4D condition, we ask the LLM to mimic an ego-centric video to further improve the diversity of the dataset. For each data generation, we prompt the LLM with the task prompt and one of the condition prompts as shown in Figure 12. The statistics TextVid are shown in Table 10. Additionally, we provide Wordcloud of TextVid in Figure 10. The examples of TextVid are shown in Appendix E.

### v2

@@ 3.4 Implementation Details
We leverage Llama2-7B, Llama2-13B [57] and Llama3-8B as the LLM backbone. Additionally, we employ the Llama-adapter [84] with an adaptation embedding length of 50. We utilize CLIP-ViT-L as the multimodal encoder. We employ a simple linear layer to project the CLIP feature into the LLM feature space. During training, the CLIP model and LLM backbone are frozen. The projection layer and additional Llama-adapter are trainable. For text-only pre-alignment, we uniformly sample the Tideos into 10 frames. We train the model on a mixture of tasks comprising Tideo summarization, Tideo QA, multi-choice Tideo QA with the ratio of 1:1:2. For zero-shot inference, we construct a memory for cross-modal projection, consisting of 2M CLIP text features sampled from captions in the TextVid dataset. TOPA-Llama2-7B and TOPA-Llama3-8B are trained on four 40G-A100 GPUs in one day. TOPA-Llama2-13B is trained in two days. More training details of TOPA and baselines are included in Appendix E.2.
@@ 4 Experiments
TOPA enables the LLM to perform various video understanding tasks as shown in Figure 2. In this section, we evaluate TOPA on multi-choice video QA and video captioning tasks. Section 4.1 evaluates TOPA on NeXT-QA [72], STAR [70], TVQA [24], recent challenging EgoSchema [39] and MVBench[29] benchmarks with the zero-shot setting. We further evaluate TOPA on multi-choice video QA with the finetuning setting (Section 4.2) and zero-shot video captioning task (Section 4.3). In Section 4.4, we conduct ablation study on the LLM prior and input video frames. We report Top-1 accuracy on multi-choice video QA benchmarks and CIDEr [58] score on video captioning benchmarks. We mainly compare TOPA with the following categories of video understanding approaches:
@@
(2) Adapting image MLLMs for video understanding [22, 81, 49]. These approaches aim to extend the image understanding capabilities of recent vision-language models (VLMs) to video understanding. Specifically, SeViLa [81] utilizes BLIP-2 for localizing and understanding key frames of a video. IG-VLM [22] converts video into a composite image by arranging the video frames into a grid layout.
(3) LLM-based video agents [42, 59, 60, 63, 82, 13, 21, 67]. This line of work leverages LLMs like GPT-3.5 and GPT-4 as an agent to understand a video by designing and executing a series of actions. The language-only agents perceive visual information via recent foundation VLMs (e.g., CLIP [48], BLIP-2 [27], LaViLa [87] and PALI [8]).
(4) Our text-only pre-alignment. Different from the above works, TOPA leverages the proposed TextVid dataset for video-LLM pre-alignment, enabling the LLM to process continuous features. Thus, it can enable performing video understanding tasks.
@@ 4.1.1 Zero-shot Results on EgoSchema
denotes results on EgoSchema subset. Results of InternVideo and FrozenBiLM are sourced from  [39]. Results of SeViLA are sourced from [45].
@@
Human Eval [39][NeurIPS 2023] - - 75.0
Gemini-1.5-Pro [50][arXiv 2024.2 ] - Gemini-1.5-Pro 63.2
@@
FrozenBiLM [77][NeurIPS 2022] - - 26.9
InternVideo [64][arXiv 2022.12] - - 32.1
LongViViT [45][CVPR  2024 ] - - 33.3
MC-ViT-L [4][ICML  2024 ] - - 44.4
InternVideo2s3-6B [66][arXiv 2024.3 ] - - 41.1
@@
SeViLA [81][NeurIPS 2023] BLIP-2 FLAN-T5-XL [10] 22.7
MVU [49][arXiv 2024.3 ] LLaVA-v1.5-13B Vicuna-13B 37.6
IG-VLM [22][arXiv 2024.3 ] LLaVA v1.6-7B Vicuna-7B 35.8
IG-VLM [22][arXiv 2024.3 ] LLaVA v1.6-13B Vicuna-13B 47.0
@@
LangRepo [21][arXiv 2024.3 ] CLIP-ViT-L Mixtral-12B [20] 41.2
Vamos [60][arXiv 2023.11] BLIP-2 Llama2-13B 36.7
Vamos [60][arXiv 2023.11] BLIP-2 GPT-3.5 41.2
Vamos [60][arXiv 2023.11] BLIP-2 GPT-4 48.3
MoReVQA [42][CVPR 2024 ] PALI-3-5B [8] PaLM-2 [2] 51.7
LLoVi [82][arXiv 2024.3 ] LaViLa GPT-3.5 50.3
VideoAgent [63][ECCV 2024 ] LaViLa GPT-4 54.1
LifelongMemory [67][arXiv 2024.3 ] LaViLa GPT-4 62.4
VideoAgent [13][ECCV 2024 ] Video-LLava [32] GPT-4 60.2
(Our Text-Only Pre-Alignment) TOPA CLIP-ViT-L Llama2-7B 41.2
TOPA CLIP-ViT-L Llama3-8B 44.2
TOPA CLIP-ViT-L Llama2-13B 51.0
Table 1 shows the results on EgoSchema full set. We compare our method against a range of recent approaches in video understanding. Our proposed text-only pre-alignment framework, despite training without real videos, shows impressive results on the EgoSchema benchmark. TOPA outperforms previous image-based adaptation approach IG-VLM and video agents LLoVi and Vamos with the same scale LLM (Llama2-7B and Llama2-13B). Moreover, TOPA shows consistent improvements when scaled up with a larger LLM backbone, indicating the effectiveness of LLMs in complex video understanding tasks.
Discussion 1: The necessity of high-quality language supervision for video understanding. Recent video pre-training approaches like LongViVit [45] and InternVideo [64], despite training on million-level web video-text data, show inferior performance on EgoSchema evaluation. These results highlight the inefficacy and inefficiency of conventional contrastive pre-training in understanding long-form videos, primarily due to noisy and simplistic language supervision. In contrast, our TOPA, trained on 721K Tideoswith high-quality language supervision, shows impressive results on EgoSchema. It indicates that, unlike image understanding which significantly benefits from leveraging web language as supervision, video understanding may require more precise and accurate language supervision to better capture the complex visual dynamics.
Discussion 2: Video agents versus end-to-end video-LLM modeling. Video agents have shown impressive results on the EgoSchema benchmark, aided by advanced LLMs and VLMs. However, a significant limitation of these approaches is their heavy reliance on the powerful LLMs. For example, the accuracy of Vamos drops by -11.6% when the GPT-4 is replaced with Llama2-13B, largely falling behind the performance of the TOPA-Llama2-13B model. The reliance on powerful closed-source LLMs restricts its application fields and introduces external overheads. Moreover, video agents make decisions based on the language format clues collected by VLMs. Converting the video content into language clues may lead to a limited upper bound compared to end-to-end modeling. Additionally, the inference speed of these approaches is another concern, since it involves multiple interactions with both VLMs and LLMs. In contrast, end-to-end video-LLM models, which condense the video into a sequence of embeddings as the input of LLM, are more efficient.
@@ 4.1.2 Zero-shot Results on NExT-QA, STAR and TVQA
4.1.3 Results on MVBench
MVBench [29] is a recent video-language understanding benchmark that covers 20 challenging video tasks, regrouped from existing video-language benchmarks. Table 3 shows the results. TOPA demonstrates impressive results compared to previous image MLLM and video MLLM. It excels particularly in tasks requiring high-level video-language understanding, such as Scene Transition (ST), Episodic Reasoning (ER), and Unexpected Action (UA). TOPA Surprisingly excels in the Action Localization (AL) task, which requires identifying the moment an action occurs. This indicates that the text-only pre-alignment enables the LLM to understand temporal visual sequences. However, TOPA struggles with tasks that demand fine-grained visual understanding, such as Moving Direction (MR), Action Antonym (AA), and Object Shuffle (OS). A common challenge in these tasks is the requirement for detailed visual understanding. For example, Action Antonym involves identifying the direction of an action, while Object Shuffle involves locating objects. TOPA struggles in these fine-grained visual tasks since it is trained with CLIP text features. The modality gap between CLIP text features and image features hinders TOPA from capturing visual details. Further video instruction tuning might address this limitation, which we leave for future work. We provide qualitative results in Appendix A.3 to illustrate TOPA’s advantages and limitations on various video understanding tasks.
@@ 4.1.3 Results on MVBench
Video MLLMs: All models take 16 frames as input
@@
Table 3: Evaluation results on MVBench. The results of other approaches are sourced from  [29]. We gray out the results of VideoChat2 since it utilizes extensive annotated downstream video data.
@@
4.2 Supervised Finetuning
In this section, we further finetune the pre-aligned TOPA models to study the benefits of TOPA for downstream supervised learning. During finetuning, TOPA directly takes the video feature as input without the cross-modal projection. More finetuning details for each dataset are provided in Appendix E.2. Table 4 shows the finetuning results on multi-choice video QA dataset. For comparison, we include baseline models without text-only pretraining. Our text-only pre-alignment consistently improves the performance across three benchmarks. Notably, TOPA-Llama2-7B achieves 67.1% accuracy on TVQA, outperforming other approaches by a large margin. These results suggest that our text-only pre-alignment, even without training on real videos, has a similar effect to conventional video-language pre-training.
@@ 4.2 Supervised Finetuning
Data-efficient finetuning. Figure 3 shows the results of finetuning LLMs with various ratios of training data. TOPA trained with 10% data achieves 64.7% Top 1 accuracy on NeXT-QA benchmark, significantly outperforming the baseline that without text-only pre-alignment. Besides, when trained with less than 20% data, the baseline model even performs worse than TOPA-zeroshot on NeXT-QA and TVQA, clearly demonstrating the effectiveness of TOPA in limited annotated data scenarios.
@@ 5 Conclusions
Acknowledgements
This work was supported by National Key R&D Program of China (No. 2023YFC3305600), the National Natural Science Foundation of China (U2336212), the Fundamental Research Funds for the Zhejiang Provincial Universities (226-2024-00208), Lu’s Graduate Education International Exchange Foundation and the National Research Foundation, Singapore under its Strategic Capability Research Centres Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore. The computational work was partially performed on resources of the National Supercomputing Centre, Singapore (https://www.nscc.sg).
@@ Acknowledgements
Appendix B: The limitations of TOPA.
Appendix C: The broader impact of TOPA.
Appendix D: The details of proposed TextVid dataset.
Appendix E.1, The details of benchmarks.
Appendix E.2, The training details of TOPA.
Appendix F: The prompts used in this paper.
Appendix G: The licenses of datasets, codes and models used in this paper.
Appendix H: Examples from TextVid.
@@ A.1 Further Discussion on Multi-Choice Video QA Task
TOPA-LLama2-7B LLM Logits 64.5 41.7 -22.8
@@
The performance gap between the EgoSchema subset and full set. Previous work [4, 45] highlights a huge performance gap between the subset and the full set of EgoSchema as shown in Table 8. While concurrent work [21, 49] introduces log-likelihood based approaches for LLM inference, which significantly improve the performance on EgoSchema subset, the issue of the performance gap still persists. In this paper, we observe that such a performance gap phenomenon also occurs in approaches based on LLM logits. However, but it diminishes or even disappears in methods employing LLM selection. We find that this phenomenon may be attributed to differences in the linguistic structures of the choices, as shown below. The choices in the subset often differ in several key works like “create”, “repair” and “clean”. The similarity or logit can effectively identify this keyword-level difference to select a more appropriate choice. Conversely, the choices in the full set display more substantial linguistic differences. These variations introduce significant language biases, i.e., some sentences naturally receive higher logits in LLM, complicating the reliance on similarity or logit for choice selection. In contrast, LLM selection methods take all the choices within the context, allowing the LLM to leverage its robust contextual understanding to select the correct choice.
@@ A.3 Qualitative Results and Analysis
We present qualitative results to illustrate the capabilities and limitations of TOPA across various video understanding tasks. Figure 4 shows qualitative results on the NExT-QA validation set. Figure 5 shows qualitative results on the EgoSchema subset. Figure 6 - 9 shows qualitative results on 20 video understanding tasks from MVBench.
@@
Appendix B Limitations.
Modality gap in CLIP. Despite the fact that TOPA achieves impressive results, a significant limitation in TOPA is the gap between the CLIP text feature and CLIP image feature. On the one hand, we use the CLIP text feature for pre-alignment, while inference is with the CLIP visual feature. The modality gap makes the performance degrade, despite employing a modality projection mechanism to mitigate it. On the other hand, the CLIP text features cannot fully capture the fine-grained visual details present in actual images, such as object locations and relationships. This limitation causes TOPA to struggle in scenarios where questions involve detailed visual information, as shown in Appendix A.3.
Appendix C Broader Impact
Appendix D The details of proposed TextVid dataset
We utilize Gemini Pro 1.0 API for our data generation process. We prompt the LLM to create textual videos along with associated annotations. To ensure a diverse dataset that covers a wide of domains, we add condition prompts including different themes, video captions, video events, and the names of main objects. Specifically, we leverage video titles from Howto100M [40], video captions from WebVid2M [3], video events from Ego4D [15], and object from Wordnet [41] as conditions to generate diverse textual videos. For Ego4D condition, we ask the LLM to mimic an ego-centric video to further improve the diversity of the dataset. Table 10 compares vocabulary sizes. Figure 11 shows that Tideos generated under different conditions have different distributions. For each data generation, we prompt the LLM with the task prompt and one of the condition prompts as shown in Figure 13. The statistics TextVid are shown in Table 10. Additionally, we provide Wordcloud of TextVid in Figure 10. The examples of TextVid are shown in Appendix H.